OpenStudy (anonymous):

g(x)=sqrt x+2 ; x=-2,-2,2,7 @jim_thompson5910

OpenStudy (anonymous):

so starting with x=-2 i would get g(-2)=sqrt-2+2 so it'd be g(-2)=sqrt0????????????

jimthompson5910 (jim_thompson5910):

the square root of 0 is _____

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

I'm assuming the function is \[\large g(x) = \sqrt{x+2}\]

jimthompson5910 (jim_thompson5910):

so (-2,0) is on the graph

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

o

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

would the domain be all real numbers too?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

there are some numbers that will not work

jimthompson5910 (jim_thompson5910):

in general, you cannot take the square root of a negative number to get a real number as an output

jimthompson5910 (jim_thompson5910):

so if x+2 is negative, then it won't work this means you have to make x+2 either 0 or positive

OpenStudy (anonymous):

x>=-2 is the domain?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

you can also see the domain when you graph (like you saw the range)

OpenStudy (anonymous):

and the range would be y>=0

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!!!!!!!!!!!!!!

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

how is the graph suppose to look?

OpenStudy (anonymous):

something like that?

jimthompson5910 (jim_thompson5910):

something like this